х P 0 .2645 1 .3518 2 .2339 os 3 2074 X 4 2065 х 5 .2208 X 0 6 or more .243976 XFor a recent period of 100 years, there were 133 major earthquakes in a particular region. Assuming that the Poisson distribu
Answer #1

See 133 Mean = 1:33 100 2 -1.33 e * 1:33 0.2645 11 P(x=0) é ^* ^ x! 0! 1 P(x=1) e-1.33 *1.33 = 0.3510 1! 2 P(x=9) - e-133 * 1